# Hidden zero modes and topology of multiband non-Hermitian systems

### v2

@@ VII Conclusions
Appendix A Other bulk-boundary correspondences
In this Appendix, we want to consider other examples which have been studied in the literature and which have lead to some misconceptions about what is and what is not topologically protected in a non-Hermitian system.
@@ A.1 Non-Hermitian SSH model
What has lead to some confusion is that there are two eigenvalues whose magnitude is going to zero with system size in a range which does not agree with the phase boundaries coming from the winding numbers. This has lead to the claim that one needs to abandon the notion of topology based on invariants for the Bloch Hamiltonian $h(k)$ [17, 18, 15]. In Ref. [17], in particular, a ’non-Bloch topological invariant’ is constructed which is meant to explain the topology of the system and the related topologically protected zero eigenmodes. This notion, however, is misguided because in part of the regime where the eigenvalues are zero there is no stability against small perturbations implying that there is no topological protection. For the specific case considered, adding small random matrices to the Hamiltonian for a finite system clearly shows that the almost zero eigenvalues are only stable in $-1/3<t_{1}<1/3$, see Fig. 4 right column. Outside of this regime, there is no topological protection in contrast to what is implied by the non-Bloch invariant constructed in Ref. [17]. On the other hand, the singular values are completely stable against these perturbations demonstrating that they are indeed topologically protected. That there are topologically protected zero eigenmodes for $-1/3<t_{1}<1/3$ can be understood in the following manner: In this case we have $\mathcal{I}_{1}=-1$ and $\mathcal{I}_{2}=+1$ as for a chiral Hermitian systems and, indeed, we can adiabatically deform the Hamiltonian (16) to a Hermitian one in this regime. This can be done, for example, by changing $\gamma/2\to-\gamma/2$ in the upper right block. Along this path, the gap never closes and the winding numbers thus remain the same as well. We conclude that the zero eigenmodes are only protected in the regime where the model is adiabatically connected to a topologically non-trivial Hermitian model. Outside of this regime, there is no protection of eigenmodes with zero energy.
@@ Appendix B Experimental detection of hidden zero modes
In this Appendix, we want to describe in more detail how hidden zero modes can be detected experimentally.
@@
An experimental system will always be finite and a hidden zero mode will be close to but not quite an eigenstate. I.e., a hidden zero mode is a long-lived metastable state. Such states will show up in various response functions—in many-body physics we are used to excitations with a finite lifetime—but the perhaps easiest way to detect them is by a Loschmidt echo type experiment. In the simplest form of such an experiment, the system is prepared in the hidden zero mode $|\Psi\rangle$ and the generalized Loschmidt echo
$\mathcal{L}(t)=\frac{\langle\Psi|\Psi(t)\rangle}{\langle\Psi(t)|\Psi(t)\rangle}$ (24)
is obtained. $\mathcal{L}(t)$ is a measure for how long the time-evolved state $|\Psi(t)\rangle$ remains close to the initial state $|\Psi\rangle$. We call Eq. (24) a generalized Loschmidt echo because the time evolution is not unitary. Note that this implies, in particular that the norm $\langle\Psi(t)|\Psi(t)\rangle$ is not conserved.
@@
Figure 5: Upper panel: $\mathcal{L}(t)$ for the Hamiltonian in example 1 of the main paper. Lower panel: $\mathcal{L}(t)$ for the SSH chain, (17). In both cases the chain has $L=70$ sites.
For example 1, the Hamiltonian has elements $h_{0}=1$, $h_{1}=2$ and the hidden zero mode is given by $v=\begin{pmatrix}\frac{\sqrt{3}}{2},&-\frac{\sqrt{3}}{4},&\dots,&\frac{\sqrt{3}}{2^{L-1}},&-\frac{\sqrt{3}}{2^{L}}\end{pmatrix}$. If we prepare the system in this state $v$, then $\mbox{Re}[\mathcal{L}(t)]\approx 1$ for times $gt<L$ where $g$ is some effective velocity which depends on the hopping amplitude. In contrast, if we prepare the system in the uniform state $\tilde{v}=\begin{pmatrix}\frac{1}{\sqrt{L}},&\frac{1}{\sqrt{L}},&\dots,&\frac{1}{\sqrt{L}},&\frac{1}{\sqrt{L}}\end{pmatrix}$ then we see a damped oscillation and the Loschmidt echo goes to zero quickly. Also shown in Fig. 5 is the case where the Hamiltonian is detuned by setting $h_{0}=2.2$ but keeping the initial state $v$. Then, there is some immediate decay clearly indicating that $v$ is not a hidden zero mode of the detuned Hamiltonian.
For the SSH model we found that one of the hidden zero modes $v$ in the regime $-1<t_{1}+\frac{\gamma}{2}<1$ is given by Eq. (21). We prepare the system in this state for $t_{1}=-0.2$ and $\gamma=4/3$. Again, the Loschmidt echo shows that this mode is indeed a long-lived metastable state. We also show $\mathcal{L}(t)$ for additional cases. One additional case is for the uniform state $\tilde{v}$ under the same Hamiltonian evolution. The other is for the hidden zero mode $v$ in the presence of a detuned Hamiltonian with the hopping amplitude changed to $t_{1}=-0.1$. In both cases the Loschmidt echo decays quickly.
We conclude that hidden zero modes are important features of non-Hermitian systems and are experimentally observable as long-lived states.

### v3

@@ VII Conclusions
Acknowledgements: The authors acknowledge support by NSERC via the Discovery grants program and by the DFG via the Research Unit FOR 2316.
Appendix A Appendix A: Other bulk-boundary correspondences
The purpose of this appendix is to consider other examples which have been studied in the literature and which have lead to some misconceptions about what is and what is not topologically protected in a non-Hermitian system.
@@ A.1 Non-Hermitian SSH model
What has lead to some confusion is that there are two eigenvalues whose magnitude is going to zero with system size in a range which does not agree with the phase boundaries coming from the winding numbers. This has lead to the claim that one needs to abandon the notion of topology based on invariants for the Bloch Hamiltonian $h(k)$ [17, 18, 15]. In Ref. [17], in particular, a ’non-Bloch topological invariant’ is constructed which is meant to explain the topology of the system and the related topologically protected zero eigenmodes. This notion, however, is misguided because in part of the regime where the eigenvalues are zero there is no stability against small perturbations implying that there is no topological protection. The theory is based on translational invariance, however, "topological features must be properties of a translationally not invariant Hamiltonian" [20]. For the specific case considered, adding small random matrices to the Hamiltonian for a finite system clearly shows that the almost zero eigenvalues are only stable in $-1/3<t_{1}<1/3$, see Fig. 4 right column. Outside of this regime, there is no topological protection in contrast to what is implied by the non-Bloch invariant constructed in Ref. [17]. On the other hand, the singular values are completely stable against these perturbations demonstrating that they are indeed topologically protected. That there are topologically protected zero eigenmodes for $-1/3<t_{1}<1/3$ can be understood in the following manner: In this case we have $\mathcal{I}_{1}=-1$ and $\mathcal{I}_{2}=+1$ as for a chiral Hermitian system and, indeed, we can adiabatically deform the Hamiltonian (16) to a Hermitian one in this regime. This can be done, for example, by changing $\gamma/2\to-\gamma/2$ in the upper right block. Along this path, the gap never closes and the winding numbers thus remain the same as well. We conclude that the zero eigenmodes are only protected in the regime where the model is adiabatically connected to a topologically non-trivial Hermitian model. Outside of this regime, there is no protection of eigenmodes with zero energy.
@@ Appendix B Experimental detection of hidden zero modes
In this appendix, we describe in more detail how hidden zero modes can be detected experimentally.
@@
An experimental system will always be finite and a hidden zero mode will be close to but not quite an eigenstate. I.e., a hidden zero mode is a long-lived metastable state. Such states will show up in various response functions—in many-body physics we are used to excitations with a finite lifetime—but perhaps the easiest way to detect them is by a Loschmidt echo type experiment. In the simplest form of such an experiment, the system is prepared in the hidden zero mode $|\Psi\rangle$ and the generalized Loschmidt echo
$\mathcal{L}(t)=\frac{\langle\Psi|\Psi(t)\rangle}{\sqrt{\langle\Psi(t)|\Psi(t)\rangle}}$ (24)
is obtained. $\mathcal{L}(t)$ is a measure for how long the time-evolved state $|\Psi(t)\rangle$ remains close to the initial state $|\Psi\rangle$. We call Eq. (24) a generalized Loschmidt echo because the time evolution is not unitary. Note that this implies, in particular, that the norm $\langle\Psi(t)|\Psi(t)\rangle$ is not conserved.
@@
Figure 5: Upper panel: $\mathcal{L}(t)$ for the Hamiltonian in example 1 of the main paper. Lower panel: $\mathcal{L}(t)$ for the SSH chain, Eq. (17). In both cases the chain has $L=70$ sites.
For example 1, the Hamiltonian has elements $h_{0}=1$, $h_{1}=2$ and the hidden zero mode is given by $v=\begin{pmatrix}\frac{\sqrt{3}}{2},&-\frac{\sqrt{3}}{4},&\dots,&\frac{\sqrt{3}}{2^{L-1}},&-\frac{\sqrt{3}}{2^{L}}\end{pmatrix}$. If we prepare the system in this state $v$, then $\mbox{Re}[\mathcal{L}(t)]\approx 1$ for times $gt<L$ where $g$ is some effective velocity which depends on the hopping amplitude. In contrast, if we prepare the system in the uniform state $\tilde{v}=\begin{pmatrix}\frac{1}{\sqrt{L}},&\frac{1}{\sqrt{L}},&\dots,&\frac{1}{\sqrt{L}},&\frac{1}{\sqrt{L}}\end{pmatrix}$ then we see a damped oscillation and the Loschmidt echo goes to zero quickly. Also shown in Fig. 5 is the case where the Hamiltonian is detuned by setting $h_{1}=2.2$ but keeping the initial state $v$. Then, there is some immediate decay clearly indicating that $v$ is not a hidden zero mode of the detuned Hamiltonian.
For the SSH model we found that one of the hidden zero modes $v$ in the regime $-1<t_{1}+\frac{\gamma}{2}<1$ is given by Eq. (21). We prepare the system in this state for $t_{1}=-0.2$ and $\gamma=4/3$. Again, the Loschmidt echo shows that this mode is indeed a long-lived metastable state. We also show again $\mathcal{L}(t)$ for additional cases. One additional case is for the uniform state $\tilde{v}$ under the same Hamiltonian evolution. The other is for the hidden zero mode $v$ in the presence of a detuned Hamiltonian with the hopping amplitude changed to $t_{1}=-0.1$. In both cases, the Loschmidt echo quickly moves away from being equal to one.
We conclude that hidden zero modes are important features of non-Hermitian systems and are experimentally observable as long-lived states. They are distinguishable from true eigenstates because the Loschmidt echo does start to decay at a time scale which increases linearly with system size $L$ whereas $\mathcal{L}(t)=1$ for all times for a true eigenstate.